# Dynamical Dark Energy Beyond Planck? Constraints from multiple CMB probes, DESI BAO and Type-Ia Supernovae

### v1

@@ Appendix A Numerical constraints on cosmological parameters and contour plots
$w_{0}$ $-0.763^{+0.066}_{-0.074}\,(-0.76^{+0.15}_{-0.13})$ $-0.765^{+0.064}_{-0.072}\,(-0.77^{+0.14}_{-0.13})$ $-0.733\pm 0.066\,(-0.73^{+0.13}_{-0.13})$
$w_{a}$ $-0.84^{+0.37}_{-0.29}\,(-0.84^{+0.63}_{-0.68})$ $-0.81^{+0.35}_{-0.28}\,(-0.81^{+0.60}_{-0.64})$ $-1.01^{+0.32}_{-0.26}\,(-1.01^{+0.55}_{-0.63})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02162\pm 0.00030\,(0.02162^{+0.00058}_{-0.00059})$ $0.02240\pm 0.00018\,(0.02240^{+0.00035}_{-0.00035})$ $0.02237\pm 0.00012\,(0.02237^{+0.00023}_{-0.00023})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1192\pm 0.0015\,(0.1192^{+0.0029}_{-0.0029})$ $0.1194\pm 0.0014\,(0.1194^{+0.0028}_{-0.0028})$ $0.11986\pm 0.00092\,(0.1199^{+0.0018}_{-0.0018})$
$\tau_{\mathrm{reio}}$ $0.064\pm 0.012\,(0.064^{+0.025}_{-0.024})$ $0.060\pm 0.012\,(0.060^{+0.024}_{-0.024})$ $0.0538\pm 0.0071\,(0.054^{+0.014}_{-0.014})$
$H_{0}$ $66.80\pm 0.68\,(66.8^{+1.4}_{-1.3})$ $67.19\pm 0.67\,(67.2^{+1.3}_{-1.3})$ $67.26\pm 0.67\,(67.3^{+1.3}_{-1.3})$
$n_{\mathrm{s}}$ $0.998\pm 0.011\,(0.998^{+0.023}_{-0.022})$ $0.9742\pm 0.0045\,(0.9742^{+0.0090}_{-0.0087})$ $0.9688\pm 0.0034\,(0.9688^{+0.0067}_{-0.0066})$
$\log(10^{10}A_{\mathrm{s}})$ $3.058\pm 0.023\,(3.058^{+0.046}_{-0.044})$ $3.062\pm 0.022\,(3.062^{+0.043}_{-0.042})$ $3.050\pm 0.013\,(3.050^{+0.027}_{-0.026})$
@@
$w_{0}$ $-0.93^{+0.12}_{-0.11}\,(-0.93^{+0.21}_{-0.24})$ $-0.92\pm 0.11\,(-0.92^{+0.22}_{-0.22})$ $-0.865\pm 0.097\,(-0.86^{+0.19}_{-0.19})$
$w_{a}$ $-0.07\pm 0.53\,(-0.07^{+0.99}_{-1.0})$ $-0.18^{+0.61}_{-0.46}\,(-0.18^{+0.99}_{-1.1})$ $-0.57^{+0.52}_{-0.44}\,(-0.57^{+0.89}_{-0.98})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02225\pm 0.00032\,(0.02225^{+0.00062}_{-0.00063})$ $0.02238\pm 0.00021\,(0.02238^{+0.00041}_{-0.00041})$ $0.02234\pm 0.00013\,(0.02234^{+0.00025}_{-0.00025})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1163\pm 0.0036\,(0.1163^{+0.0071}_{-0.0071})$ $0.1172\pm 0.0028\,(0.1172^{+0.0054}_{-0.0054})$ $0.1201\pm 0.0013\,(0.1201^{+0.0025}_{-0.0025})$
$\tau_{\mathrm{reio}}$ $0.059\pm 0.014\,(0.059^{+0.028}_{-0.028})$ $0.058\pm 0.014\,(0.058^{+0.028}_{-0.029})$ $0.0518\pm 0.0078\,(0.052^{+0.016}_{-0.015})$
$H_{0}$ $66.7\pm 1.6\,(66.7^{+3.1}_{-3.3})$ $66.9\pm 1.5\,(66.9^{+2.9}_{-2.9})$ $67.6\pm 1.2\,(67.6^{+2.3}_{-2.3})$
$n_{\mathrm{s}}$ $0.972\pm 0.016\,(0.972^{+0.032}_{-0.032})$ $0.9678\pm 0.0067\,(0.968^{+0.013}_{-0.013})$ $0.9652\pm 0.0040\,(0.9652^{+0.0079}_{-0.0078})$
$\log(10^{10}A_{\mathrm{s}})$ $3.047\pm 0.030\,(3.047^{+0.059}_{-0.059})$ $3.044\pm 0.028\,(3.044^{+0.054}_{-0.056})$ $3.040\pm 0.016\,(3.040^{+0.031}_{-0.031})$
@@
$w_{0}$ $-0.79^{+0.13}_{-0.11}\,(-0.79^{+0.22}_{-0.25})$ $-0.80^{+0.12}_{-0.11}\,(-0.80^{+0.22}_{-0.24})$ $-0.74\pm 0.11\,(-0.74^{+0.21}_{-0.21})$
$w_{a}$ $-0.63\pm 0.58\,(-0.6^{+1.1}_{-1.1})$ $-0.61\pm 0.58\,(-0.6^{+1.1}_{-1.1})$ $-0.995\pm 0.54\,(-1.0^{+1.0}_{-1.1})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02226\pm 0.00032\,(0.02226^{+0.00062}_{-0.00062})$ $0.02239\pm 0.00021\,(0.02239^{+0.00042}_{-0.00041})$ $0.02235\pm 0.00013\,(0.02235^{+0.00026}_{-0.00025})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1166\pm 0.0038\,(0.1166^{+0.0075}_{-0.0075})$ $0.1170\pm 0.0027\,(0.1170^{+0.0052}_{-0.0053})$ $0.1201\pm 0.0013\,(0.1201^{+0.0025}_{-0.0024})$
$\tau_{\mathrm{reio}}$ $0.058\pm 0.015\,(0.058^{+0.029}_{-0.029})$ $0.057\pm 0.014\,(0.057^{+0.027}_{-0.028})$ $0.0514\pm 0.0077\,(0.051^{+0.015}_{-0.015})$
$H_{0}$ $66.8^{+1.4}_{-1.2}\,(66.8^{+2.5}_{-2.8})$ $66.8^{+1.4}_{-1.1}\,(66.8^{+2.4}_{-2.7})$ $67.2^{+1.1}_{-0.97}\,(67.2^{+2.0}_{-2.1})$
$n_{\mathrm{s}}$ $0.972\pm 0.017\,(0.972^{+0.033}_{-0.032})$ $0.9682\pm 0.0067\,(0.968^{+0.013}_{-0.013})$ $0.9654\pm 0.0040\,(0.9654^{+0.0078}_{-0.0079})$
$\log(10^{10}A_{\mathrm{s}})$ $3.045\pm 0.030\,(3.045^{+0.058}_{-0.060})$ $3.043\pm 0.027\,(3.043^{+0.053}_{-0.054})$ $3.039\pm 0.015\,(3.039^{+0.030}_{-0.031})$

### v2

@@ Appendix A Numerical constraints on cosmological parameters and contour plots
$w_{0}$ $-0.764\pm 0.058\,(-0.76^{+0.12}_{-0.11})$ $-0.774\pm 0.057\,(-0.77^{+0.11}_{-0.11})$ $-0.759\pm 0.057\,(-0.76^{+0.11}_{-0.11})$
$w_{a}$ $-0.78^{+0.26}_{-0.23}\,(-0.78^{+0.46}_{-0.50})$ $-0.71\pm 0.23\,(-0.71^{+0.44}_{-0.47})$ $-0.79^{+0.23}_{-0.20}\,(-0.79^{+0.40}_{-0.45})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02259\pm 0.00016\,(0.02259^{+0.00032}_{-0.00032})$ $0.02265\pm 0.00012\,(0.02265^{+0.00024}_{-0.00023})$ $0.02251\pm 0.00011\,(0.02251^{+0.00021}_{-0.00021})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1192\pm 0.0012\,(0.1192^{+0.0022}_{-0.0024})$ $0.1188\pm 0.0011\,(0.1188^{+0.0022}_{-0.0022})$ $0.11895\pm 0.00087\,(0.1189^{+0.0017}_{-0.0018})$
$\tau_{\mathrm{reio}}$ $0.068\pm 0.011\,(0.068^{+0.023}_{-0.022})$ $0.067\pm 0.011\,(0.067^{+0.022}_{-0.022})$ $0.0554\pm 0.0072\,(0.055^{+0.014}_{-0.014})$
$H_{0}$ $66.92\pm 0.55\,(66.9^{+1.1}_{-1.1})$ $66.94\pm 0.55\,(66.9^{+1.1}_{-1.1})$ $66.85\pm 0.57\,(66.9^{+1.1}_{-1.1})$
$n_{\mathrm{s}}$ $0.9759\pm 0.0068\,(0.976^{+0.014}_{-0.013})$ $0.9713\pm 0.0040\,(0.9713^{+0.0080}_{-0.0077})$ $0.9720\pm 0.0033\,(0.9720^{+0.0064}_{-0.0064})$
$\log(10^{10}A_{\mathrm{s}})$ $3.066\pm 0.021\,(3.066^{+0.040}_{-0.040})$ $3.070\pm 0.021\,(3.070^{+0.040}_{-0.041})$ $3.045\pm 0.013\,(3.045^{+0.026}_{-0.025})$
@@
$w$ $-0.888\pm 0.057\,(-0.89^{+0.11}_{-0.11})$ $-0.882\pm 0.056\,(-0.88^{+0.11}_{-0.11})$ $-0.860\pm 0.056\,(-0.86^{+0.11}_{-0.11})$
$w_{a}$ $-0.27^{+0.26}_{-0.23}\,(-0.27^{+0.46}_{-0.50})$ $-0.29^{+0.25}_{-0.22}\,(-0.29^{+0.43}_{-0.48})$ $-0.46^{+0.23}_{-0.21}\,(-0.46^{+0.42}_{-0.45})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02226\pm 0.00032\,(0.02226^{+0.00063}_{-0.00062})$ $0.02241\pm 0.00020\,(0.02241^{+0.00040}_{-0.00039})$ $0.02232\pm 0.00016\,(0.02232^{+0.00031}_{-0.00030})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1161^{+0.0018}_{-0.0016}\,(0.1161^{+0.0033}_{-0.0036})$ $0.1163^{+0.0017}_{-0.0015}\,(0.1163^{+0.0030}_{-0.0033})$ $0.1178\pm 0.0012\,(0.1178^{+0.0023}_{-0.0024})$
$\tau_{\mathrm{reio}}$ $0.059\pm 0.014\,(0.059^{+0.027}_{-0.027})$ $0.058\pm 0.014\,(0.058^{+0.027}_{-0.027})$ $0.0481^{+0.0085}_{-0.0076}\,(0.048^{+0.017}_{-0.016})$
$H_{0}$ $67.25\pm 0.63\,(67.3^{+1.2}_{-1.2})$ $67.32\pm 0.61\,(67.3^{+1.2}_{-1.2})$ $67.43\pm 0.60\,(67.4^{+1.2}_{-1.2})$
$n_{\mathrm{s}}$ $0.973\pm 0.015\,(0.973^{+0.030}_{-0.029})$ $0.9690\pm 0.0056\,(0.969^{+0.011}_{-0.011})$ $0.9691\pm 0.0052\,(0.969^{+0.010}_{-0.010})$
$\log(10^{10}A_{\mathrm{s}})$ $3.046\pm 0.030\,(3.046^{+0.058}_{-0.058})$ $3.043\pm 0.027\,(3.043^{+0.053}_{-0.053})$ $3.025^{+0.018}_{-0.016}\,(3.025^{+0.034}_{-0.034})$
@@
$w_{0}$ $-0.799\pm 0.061\,(-0.80^{+0.12}_{-0.12})$ $-0.797\pm 0.061\,(-0.80^{+0.12}_{-0.11})$ $-0.774\pm 0.059\,(-0.77^{+0.12}_{-0.11})$
$w_{a}$ $-0.55\pm 0.27\,(-0.55^{+0.51}_{-0.54})$ $-0.56^{+0.28}_{-0.24}\,(-0.56^{+0.49}_{-0.54})$ $-0.71^{+0.26}_{-0.22}\,(-0.71^{+0.46}_{-0.51})$
$\Omega_{\mathrm{b}}h^{2}$ $0.02225\pm 0.00032\,(0.02225^{+0.00062}_{-0.00061})$ $0.02239\pm 0.00020\,(0.02239^{+0.00040}_{-0.00040})$ $0.02229\pm 0.00015\,(0.02229^{+0.00031}_{-0.00030})$
$\Omega_{\mathrm{c}}h^{2}$ $0.1169^{+0.0017}_{-0.0016}\,(0.1169^{+0.0031}_{-0.0034})$ $0.1169^{+0.0016}_{-0.0014}\,(0.1169^{+0.0028}_{-0.0031})$ $0.1181\pm 0.0012\,(0.1181^{+0.0022}_{-0.0024})$
$\tau_{\mathrm{reio}}$ $0.058\pm 0.014\,(0.058^{+0.028}_{-0.028})$ $0.057\pm 0.014\,(0.057^{+0.027}_{-0.027})$ $0.0477^{+0.0086}_{-0.0074}\,(0.048^{+0.015}_{-0.018})$
$H_{0}$ $66.56\pm 0.59\,(66.6^{+1.2}_{-1.2})$ $66.63\pm 0.58\,(66.6^{+1.1}_{-1.1})$ $66.66\pm 0.57\,(66.7^{+1.1}_{-1.1})$
$n_{\mathrm{s}}$ $0.972\pm 0.015\,(0.972^{+0.030}_{-0.029})$ $0.9681\pm 0.0055\,(0.968^{+0.011}_{-0.011})$ $0.9682\pm 0.0052\,(0.968^{+0.010}_{-0.010})$
$\log(10^{10}A_{\mathrm{s}})$ $3.045\pm 0.030\,(3.045^{+0.060}_{-0.059})$ $3.042\pm 0.027\,(3.042^{+0.053}_{-0.054})$ $3.025^{+0.018}_{-0.016}\,(3.025^{+0.032}_{-0.036})$
@@
Appendix B The role of the reionization optical depth and matter density parameters
As I showed in the main text of this paper, removing Planck large-scale temperature and polarization data – and above all E-mode polarization measurements – or considering CMB experiments different from Planck significantly weakens the preference for dynamical dark energy. As I pointed out in the main manuscript, this is due to at least two different effects. The first is related to the ISW contribution, which mainly affects large-scale temperature modes. Removing the TT data slightly weakens the DDE signal. However, I concluded that the most significant effect comes from removing large-scale E-mode polarization measurements.
As we originally noted in Ref. Giarè et al. (2024c), Planck $-\ell$ E-mode polarization data typically favor lower values of the reionization optical depth $\tau$, which correlates with several parameters that are often primarily involved in anomalous cosmological results. For example, after the second data release of DESI BAO measurements, a few independent groups pointed out that such low values of $\tau$ might play a primary role in driving the total neutrino mass towards exceedingly small values Sailer et al. (2025); Jhaveri et al. (2025).
Here I argue that large-scale constraints on $\tau$ trigger a domino effect of parameter correlations that plays a key role in strengthening the evidence for DDE. In particular, due to the well-known degeneracy between $\tau$ and the scalar amplitude $A_{s}$ – CMB anisotropies are primarily sensitive to the combination $A_{s}e^{-2\tau}$ – a lower $\tau$ leads to a lower inferred $A_{s}$. This, in turn, impacts the lensing-induced smoothing of the CMB peaks, which depends on the interplay among $(A_{s},n_{s},\Omega_{m})$. To match the observed lensing amplitude, a smaller $A_{s}$ is typically compensated by a larger $n_{s}$ and a smaller $\Omega_{m}$. Since in several dataset combinations the evidence for dynamical dark energy becomes stronger at higher values of $\Omega_{m}$, this sequence of adjustments can indirectly weaken the DDE signal.
Figure 8: Two-dimensional marginalized contours at 68% and 95% CL in the ($w_{0}$, $w_{a}$) plane for datasets involving various CMB experiments other than Planck, with different priors on $\tau$ indicated in the legend. The black dashed lines correspond to the standard $\Lambda$CDM values, $w_{0}=-1$ and $w_{a}=0$.
To support this argument, I perform two additional analyses:
  1. 1.
    When studying constraints from CMB experiments other than Planck, I replace the conservative prior on the reionization optical depth adopted in the main analysis ($\tau=0.065\pm 0.015$) with the Planck-driven prior $\tau=0.055\pm 0.007$, and check how the results for combinations such as WMAP+ACT and WMAP+SPT change compared to the baseline results discussed in the main text. If the argument is true, a lower value of $\tau$ is expected to strengthen the preference for DDE
  2. 2.
    When I remove large-scale polarization measurements in combinations that involve Planck CMB data, this automatically leads to larger uncertainties on $\tau$ (with a tendency toward higher values) and $\Omega_{m}$. To test whether the preference for DDE is weakened by lower values of $\Omega_{m}$ values, I impose a prior $\Omega_{m}=0.29\pm 0.01$ and check how the posterior shifts.
The results of the first test are shown in Fig. 8. As one can see, in combinations that involve ACT+WMAP data, the preference for DDE is significantly increased when adopting a Planck-driven prior on $\tau$. For example, when considering ACT+WMAP+DESI+PP with a prior $\tau=0.055\pm 0.007$, I obtain $w=-0.844\pm 0.055$ and $w_{a}=-0.57^{+0.22}_{-0.19}$ (both at 68% CL). As seen from the figure and by direct comparison with Tab. 2, this preference is significantly stronger than in the results presented in the main text. Similarly, replacing PP with DESy5 gives $w=-0.759\pm 0.056$ and $w_{a}=-0.80^{+0.23}_{-0.20}$, further strengthening the preference for DDE already present in this combination. That said, the impact of a different $\tau$ prior is much more modest in combinations involving SPT+WMAP. In these cases the results barely change, see also Fig. 8. Specifically, imposing a prior $\tau=0.055\pm 0.007$ for WMAP+SPT+DESI+PP yields $w=-0.883\pm 0.056$ and $w_{a}=-0.30^{+0.24}_{-0.22}$, while for WMAP+SPT+DESI+DESy5 I obtain $w=-0.796\pm 0.059$ and $w_{a}=-0.57^{+0.27}_{-0.24}$. The larger uncertainties of SPT mitigate this effect, leading to results similar to those discussed in the main text with the more conservative prior on $\tau$.
Figure 9: Two-dimensional marginalized contours at 68% and 95% CL in the ($w_{0}$, $w_{a}$) plane for different combinations of the Planck likelihoods without large-scale temperature and polarization measurements, with and without imposing a prior on $\Omega_{m}$. The black dashed lines represent the standard $\Lambda$CDM values, $w_{0}=-1$ and $w_{a}=0$.
The results of the second test are shown in Fig. 9. As seen from the figure, assuming this prior on $\Omega_{m}$ in combinations of data that involve Planck CMB measurements without large-scale temperature and polarization further pushes the contours towards a cosmological constant, decreasing even more the strength of the evidence for DDE. Specifically, imposing a prior $\Omega_{m}=0.29\pm 0.01$ for Planck(no low-L)+DESI+PP gives $w=-0.895\pm 0.053$ and $w_{a}=-0.34^{+0.23}_{-0.20}$, while replacing PP with DESy5 for Planck(no low-L)+DESI+DESy5 yields $w=-0.825\pm 0.054$ and $w_{a}=-0.55^{+0.25}_{-0.21}$.
These results confirm the key role played by E-mode polarization measurements and the correlations among these quantities, further clarifying the nature of the differences between different subsets of CMB data and the importance of CMB in the analysis.